OpenStudy (jazzkakes):

What's the difference between a multiple and a least common multiple??? Please help me !!!....

OpenStudy (jackthegreatest):



OpenStudy (jackthegreatest):



OpenStudy (loser66):

???

OpenStudy (jackthegreatest):

@Loser66 ??????

OpenStudy (loser66):

Suppose you have 2 numbers : 3 and 9 Multiple them is 3*9 =27 The least common multiple of 3 and 9 is 9 why? because 9 divided by both 3 and 9

OpenStudy (loser66):

That is what the LEAST COMMON MULTIPLE means

OpenStudy (loser66):

I made the question marks on your respond because I don't get what it means.

OpenStudy (jackthegreatest):

oh well, its not wrong

OpenStudy (welshfella):

yes 27 , 18 and 9 are all common multiples of 3 and 9 but the least common miutiple is 9.

OpenStudy (sshayer):

2,3,4 L.C.M=4*3=12 multiples=12,24,36,....